# ReCo-Diff: Explore Retinex-Based Condition Strategy in Diffusion Model for Low-Light Image Enhancement

### v1

@@ 4.3 Qualitative Evaluation
SMG [44] CVPR’23 25.852 0.869 0.248 41.647
FourLLIE [24] MM’23 22.462 0.814 0.296 61.380
UHDFour [24] ICLR’23 26.226 0.900 0.239 39.956
DiffLL [14] SIGGRAPH Asia’23 24.330 0.843 0.245 49.098
PyDiff [57] IJCAI’23 25.753 0.897 0.159 36.263
ReCo-Diff (Ours) 27.347 0.912 0.121 27.233
@@
The qualitative evaluations on LOL and LOL-v2 are shown in Figs. 1, 4 and 5. As indicated by the visual comparisons, our ReCo-Diff shows superior enhancement capability and generates images with more pleasing perceptual quality. Specifically, in the first row of Fig. 4, previous methods fail to reconstruct the detailed textures of the blanket, while our ReCo-Diff provides rich details and mitigates the color gap. As for images of the middle row, although most methods enhance the white regions well, they produce incorrect color and artifacts in the regions with complex scenes. Notably, our method not only restores the print on the carton but preserves the color consistency. Furthermore, the bottom row exhibits that our ReCo-Diff is capable of recovering the vulnerable content vanishing in the results of other methods, which reasonably indicates the superior capacity of restoring naturalistic details. In Fig. 5, although PyDiff [57] shows a similar overall look to ours, the ReCo-Diff enhances the green cup and the red text with correct color and better saturation. Hence, our ReCo-Diff yields more visually pleasing results as compared to baselines, supporting our method’s excellent performance in quantitative evaluation. More results are provided in supplementary material.
4.4 Ablation Study
@@
Figure 6: Visual comparison on the LOL/LOL-v2 datasets for investigating the contribution of key techniques of our ReCo-Diff.
@@
Contribution of ANet, RNet, and Conditional UNet. As shown in Tab. 4, we conduct experiments of several ablation settings by removing different components from the framework individually. The “w/o UNet-Cond" denotes the removal of conditional layers in the UNet. Compared with all ablation settings, our full setting yields the best performance. In the case of the removal of FRCM leads to the decrease of PSNR value by an average of 0.364dB below the baseline, which demonstrates the effectiveness of introducing the multi-scale features from ANet into the refinement process. The comparison between “w/o UNet-Cond" and the baseline shows a similar degradation level, resulting in an average reduction of 0.349dB, which proves the necessity of integrating Retinex features into the denoising network. Notably, by comparing “w/o IRCM" and “w/ ALL", we observe a significant decline that surpasses the aforementioned two cases, which serves to highlight the crucial role of our IRCM owing to its well-designed residual refinement manner. Furthermore, the comparison between all cases under “w/ ANet" and “w/o ANet" exhibits an average drop of 0.272dB on PSNR, illustrating that the better Retinex-based condition provided by ANet eventually improves the visual quality of generated images.
@@
Additionally, results in Fig. 6 depict that “w/o FRCM" and “UNet-Cond" lead to the reduction of regions with detailed textures, such as uneven surfaces and closely arranged chairs. Although “w/o IRCM" preserves details better, the lack of direct guidance of Retinex components induces the color shift and unnatural illumination. Fig. 7 shows that ANet reduces the noise (top rows) and enhances the illumination (bottom rows) in early steps, and improves the quality of Retinex-based conditions in the whole process.
@@
Table 4: Ablation studies on LOL for investigating the contribution of key techniques both in the learning and refinement stages.
@@
Learning Refinement PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
w/ ANet w/ ALL 27.626 0.884 0.090
w/o FRCM 27.344 0.875 0.112
w/o IRCM 27.262 0.872 0.119
w/o UNet-Cond 27.278 0.875 0.113
w/o ANet w/ ALL 27.301 0.877 0.103
w/o FRCM 26.956 0.869 0.131
w/o IRCM 26.783 0.864 0.140
w/o UNet-Cond 26.967 0.868 0.134
Table 5: Ablation studies of training settings of ANet for investigating the effects of the joint loss function.
Settings LOL LOL-v2-real
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
w/ Both 27.626 0.884 0.090 29.306 0.906 0.105
w/o SSIM 27.502 0.875 0.101 29.163 0.891 0.118
w/o Hist 27.508 0.875 0.102 29.156 0.892 0.119
w/o Both 27.446 0.872 0.106 29.061 0.885 0.124
@@
Figure 7: Ablation study of ANet in terms of reflectance (row 1,3) and illumination (row 2,4) maps. Images in orange box shows the effects of ANet corresponding to Tab. 4, and green box shows the effects of SSIM and Hist loss corresponding to Tab. 5.
Joint Losses of ANet. Tab. 5 lists the results of different settings for training ANet. The “w/ Both" and “w/o Both" denote that using the original $\mathcal{L}_{ja}$ and $\mathcal{L}_{je}$ and removing both SSIM loss and Hist loss, respectively. First, “w/ Both" reasonably achieves the best performance compared with all ablation settings. The settings of “w/o SSIM" and “w/o Hist" show degradation on all metrics, demonstrating the effects of SSIM loss and Hist loss respectively. Comparing “w/ Both" and “w/o Both" shows a favorable gain of 0.210dB and illustrates the necessity of using the constraint of both the structure and histogram. Furthermore, we can see the effects of SSIM loss and Hist loss by comparing Retinex output at step 8 in Fig. 7. The reflectance and illumination maps in the green box contain more artifacts and unsatisfactory brightness distribution. More ablation studies and analysis are provided in supplementary material.
5 Conclusion
@@
In this paper, we propose an LLIE-specific diffusion model with a Retinex-based condition strategy, ReCo-Diff, for low-light image enhancement. Our ReCo-Diff combines Retinex theory and diffusion model in two stages: Retinex-based condition learning and conditional refinement. In the learning stage, we utilize DNet to obtain the initial priors and then provide robust Retinex-based conditions by ANet. In the refinement stage, the prepared conditions are integrated into both UNet and RNet through conditional layers and RCMs to control the diffusion process and produce better color and details. Extensive experiments demonstrate that our ReCo-Diff outperforms state-of-the-art methods on representative benchmarks. We also hope to encourage more attention to develop DMs incorporating intrinsic characteristics of low-level vision.

### v2

@@ 4.1. Experimental Settings
SMG (Xu et al., 2023) CVPR’23 25.852 0.869 0.248 41.647
FourLLIE (Li et al., 2023a) MM’23 22.462 0.814 0.296 61.380
UHDFour (Li et al., 2023a) ICLR’23 26.226 0.900 0.239 39.956
DiffLL (Jiang et al., 2023) SIGGRAPH Asia’23 24.330 0.843 0.245 49.098
PyDiff (Zhou et al., 2023) IJCAI’23 25.753 0.897 0.159 36.263
JoReS-Diff (Ours) 27.347 0.912 0.121 27.233
Table 3. Quantitative comparison on the ISTD (Wang et al., 2018).
Methods ISTD
PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$
Shadow Image 20.56 0.908 10.86
DC-ShadowNet (Jin et al., 2021) ICCV’21 26.38 0.917 6.62
EMNet (Zhu et al., 2022b) AAAI’22 29.98 0.940 5.28
BMNet (Zhu et al., 2022a) CVPR’22 30.26 0.957 5.06
ShadowFormer (Guo et al., 2023a) AAAI’23 30.47 0.958 4.79
LFG-Diff (Mei et al., 2024) WACV’24 30.64 0.963 4.93
JoReS-Diff (Ours) 30.89 0.971 4.66
Table 4. Quantitative comparison on MIT-Adobe FiveK (Vladimir et al., 2011).
Methods MIT-Adobe FiveK
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
RetinexNet (Wei et al., 2018) BMVC’18 12.515 0.671 0.254
KinD (Zhang et al., 2019) MM’19 16.203 0.784 0.150
Uformer (Wang et al., 2022) CVPR’22 21.917 0.871 0.085
Restormer (Zamir et al., 2022a) CVPR’22 24.923 0.911 0.058
LLFormer (Wang et al., 2023c) AAAI’23 25.752 0.923 0.045
JoReS-Diff (Ours) 25.669 0.929 0.042
Implementation Details. Our method is trained using an NVIDIA RTX A100 GPU with 600k iterations. We use Adam optimizer with the momentum as $(0.9,0.999)$. The initial learning rate is set to $1\!\times\!10^{-4}$ and decays by a factor of 0.5 after every $1\!\times\!10^{5}$ iterations. The batch size and patch size are set to 8 and 256$\times$256. We use the similar architecture of (Saharia et al., 2022b) as our denoising U-Net $\bm{\epsilon}_{\theta}$. In the inference stage, we follow the DDIM (Song et al., 2020) and output by 8 steps.
Datasets. We evaluate our method on several datasets. The LOL dataset (Wei et al., 2018) includes 485 low/normal-light training pairs and 15 testing pairs. The LOL-v2 dataset (Yang et al., 2021) has two parts, the real part is an extension of LOL with 689/100 pairs and the synthetic part includes 900/100 pairs for training and testing. The UHD-LL dataset (Li et al., 2023a) includes 2000 training and 150 testing pairs. The ISTD dataset (Wang et al., 2018) includes 1330 training and 540 testing triplets. The MIT-Adobe FiveK dataset (Vladimir et al., 2011) includes 4500 training and 500 testing pairs.
Metrics. We employ peak signal-to-noise ratio (PSNR), structural similarity (SSIM) (Wang et al., 2004), perceptual image patch similarity (LPIPS) (Zhang et al., 2018), and Fréchet Inception Distance (FID) (Heusel et al., 2017) for evaluation.
Compared Methods. We compare with a rich collection of state-of-the-art LLIE methods, including LIME (Guo et al., 2016), RetinexNet (Wei et al., 2018), KinD (Zhang et al., 2019), DRBN (Yang et al., 2020), Zero-DCE (Guo et al., 2020), EnlightGAN (Jiang et al., 2021), MIRNet (Zamir et al., 2022b), SNR (Xu et al., 2022), PairLIE (Fu et al., 2023), SMG (Xu et al., 2023), FourLLIE (Li et al., 2023a), Retinexformer (Cai et al., 2023), UHDFour (Li et al., 2023a) and diffusion-based LLIE methods including Diff-Retinex (Yi et al., 2023), DiffLL (Jiang et al., 2023), and PyDiff (Zhou et al., 2023). For the shadow removal and exposure adjustment tasks, we select DC-ShadowNet (Jin et al., 2021), EMNet (Zhu et al., 2022b), BMNet (Zhu et al., 2022a), ShadowFormer (Guo et al., 2023a), LFG-Diff (Mei et al., 2024), Uformer (Wang et al., 2022), Restormer (Zamir et al., 2022a), LLFormer (Wang et al., 2023c).
4.2. Quantitative Evaluation
Tables 1 and 2 show the comparisons on LOL, LOL-v2, and UHD-LL. It is clear that our JoReS-Diff achieves consistent and significant performance gain over all competing methods. Specifically, our method provides significant improvement of 0.155 dB/0.216 dB on LOL/LOL-v2-real datasets respectively, establishing the new state-of-the-art with PSNR values of 26.491 dB/24.836 dB. Furthermore, our method achieves similar performance in terms of SSIM, yielding the best values of 0.876/0.897. On the LOL-v2-synthetic test set, our JoReS-Diff improves the PSNR by 0.042 than the second-best method. On UHD-LL, we also achieve considerable performance on PSNR and SSIM of 1.121 dB and 0.012, proving our capability of generalization. Although our JoReS-Diff obtains several second-best LPIPS values, it outperforms other diffusion-based LLIE methods and achieves the best PSNR, SSIM and FID on each case, which indicates the overall performance on perceptual metrics is still competitive. Consequently, the considerable performance shows the capability of suppressing noise and preserving color and details by Retinex and semantic priors, which demonstrates the effectiveness of our proposed joint condition strategy in resolving LLIE task.
As shown in Tables 3 and 4, we report the results on ISTD (Wang et al., 2018) and MIT-Adobe-FiveK (Vladimir et al., 2011) to evaluate the capability of other image enhancement tasks, i.e. shadow removal and exposure adjustment. We can observe the remarkable performance of our JoReS-Diff from the comparisons. To be specific, our method improves the PSNR/SSIM by 0.25 dB/0.008 than the latest method LFG-Diff, demonstrating the effectiveness of the joint prior in the shadow removal task. On FiveK (Vladimir et al., 2011), JoReS still outperforms previous strong baselines, such as Uformer and Restormer, and provides competitive performance comparing with LLFormer. The slight drop of PSNR may cause by the gap between the PASCAL-Context and FiveK.
4.3. Qualitative Evaluation
Figure 5. Visual comparison of our JoReS-Diff and the compared LLIE methods on the LOL dataset.
Figure 6. Visual comparison of our JoReS-Diff and the compared LLIE methods on the UHD-LL dataset.
@@ 4.3. Qualitative Evaluation
The qualitative evaluations on LOL and LOL-v2 are shown in Figs. 1, 5 and 6. As indicated by the visual comparisons, our JoReS-Diff shows superior enhancement capability and generates images with more pleasing perceptual quality. Specifically, in the first row of Fig. 5, previous methods fail to reconstruct the detailed textures of the blanket, while our JoReS-Diff provides rich details and mitigates the color gap. As for images of the middle row, although most methods enhance the white regions well, they produce incorrect color and artifacts in the regions with complex scenes. Notably, our method not only restores the print on the carton but preserves the color consistency. Furthermore, the bottom row exhibits that our JoReS-Diff is capable of recovering the vulnerable content vanishing in the results of other methods, which reasonably indicates the superior capacity of restoring naturalistic details. In Fig. 6, although PyDiff (Zhou et al., 2023) shows a similar overall look to ours, the JoReS-Diff enhances the green cup and the red text with correct color and better saturation. Hence, our JoReS-Diff yields more visually pleasing results as compared to baselines, supporting our method’s excellent performance in quantitative evaluation.
4.4. Ablation Study
@@
Figure 7. Visual comparison on the LOL/LOL-v2 datasets for investigating the contribution of key techniques of our JoReS-Diff.
@@
Contribution of Retinex prior incorporation. As shown in Table 5, we conduct experiments of several ablation settings by removing different components from the framework individually. The “w/o UNet-Cond” denotes the removal of Retinex attention layers in the UNet. Compared with all ablation settings, our full setting yields the best performance. In the case of the removal of FRCM leads to the decrease of PSNR value by an average of 0.565 dB below the baseline, which demonstrates the effectiveness of introducing the multi-scale features from ANet into the refinement process. The comparison between “w/o UNet-Cond” and the baseline shows a similar degradation level, resulting in an average reduction of 0.676 dB, which proves the necessity of integrating Retinex features into UNet. Notably, by comparing “w/o IRCM” and “w/ ALL”, we observe a significant decline (0.864 dB) that surpasses the aforementioned two cases, which serves to highlight the crucial role of our IRCM owing to its well-designed residual refinement manner. Furthermore, the comparison between all cases under “w/ ANet” and “w/o ANet” exhibits an average drop of 0.568 dB on PSNR, illustrating that the better Retinex-based condition provided by ANet eventually improves the visual quality of generated images. Under the case of “w/ ANet”, the settings of “w/o SSIM loss” and “w/o Hist loss” show degradation on all metrics, demonstrating the effects of SSIM loss and Hist loss respectively. Comparing “w/ ALL” and “w/o SSIM & Hist losses” shows a favorable gain of 0.459 dB and illustrates the necessity of using the constraint of both the structure and histogram. Notably, directly multiplying the outputs of ANet produce unfavorable results since the model capacity is similar with RetinexNet, while we can still obtain favorable results by using them as condition.
@@
Additionally, results in Fig. 7 depict that “w/o FRCM” and “UNet-Cond” lead to the reduction of regions with detailed textures, such as uneven surfaces and closely arranged chairs. Although “w/o IRCM” preserves details better, the lack of direct guidance of Retinex components induces the color shift and unnatural illumination. Fig. 8 shows that ANet reduces the noise (top rows) and enhances the illumination (bottom rows) in early steps, and improves the quality of Retinex-based conditions in the whole process. Furthermore, we can see the effects of SSIM and Hist losses by comparing Retinex output at step 8. The reflectance and illumination maps in the green box contain more artifacts and unsatisfactory brightness distribution.
@@
Table 5. Ablation studies on LOL for investigating the contribution of key techniques of Retinex prior incorporation.
@@
Learning Refinement PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Runtime(s)
w/ ANet w/ ALL 26.491 0.876 0.092 0.478
w/o FRCM 25.834 0.864 0.133 0.458
w/o IRCM 25.487 0.859 0.145 0.421
w/o UNet-Cond 25.673 0.861 0.135 0.414
w/o SSIM loss 26.202 0.866 0.106 -
w/o Hist loss 26.194 0.866 0.105 -
w/o SSIM & Hist 26.032 0.864 0.108 -
w/o ANet w/ ALL 25.736 0.862 0.116 0.446
w/o FRCM 25.264 0.854 0.149 -
w/o IRCM 25.012 0.849 0.168 -
w/o UNet-Cond 25.202 0.853 0.154 -
ANet Output Multiplication 17.129 0.675 0.397 -
Table 6. Ablation studies on LOL for investigating the contribution of key techniques of semantic prior incorporation.
Settings PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Runtime(s)
w/ ALL 26.491 0.876 0.092 0.478
w/o CA 26.183 0.865 0.131 0.472
w/o SA 26.024 0.863 0.118 0.472
w/o CA & SA 25.901 0.858 0.143 0.466
Contribution of semantic prior incorporation. As shown in Table 6, we conduct the ablation studies to investigate the contribution of cross-attention and self-attention in the semantic attention layers. By comparing the performance degradation caused by “w/o CA” and “w/o SA”, it is conspicuous that the self-attention mainly benefits on PSNR and SSIM, while the cross-attention can improve LPIPS more. The results reasonably support the motivation of designing the semantic attention layers in Section 3.3.
@@
Figure 8. Ablation study of ANet in terms of reflectance (row 1,3) and illumination (row 2,4) maps. Images in orange box shows the effects of ANet, and green box shows the effects of SSIM and Hist loss corresponding to Table 5.
5. Conclusion
@@
In this paper, we propose a diffusion model with joint Retinex and semnatic priors, JoReS-Diff, for image enhancement tasks. JoReS-Diff combines Retinex prior and diffusion model in two stages: Retinex-based condition learning and conditional refinement. In the learning stage, we utilize DNet to obtain initial decomposed maps and then provide robust Retinex prior by ANet. In the refinement stage, Retinex prior is integrated by Retinex attention layers and RCMs in RNet to control the diffusion process and produce better color and details. The semantic prior is provided by an off-the-shelf segmentation model and incorporated by semantic attention layers with both self- and cross-attention, preserving semantic consistency of the output. Extensive experiments demonstrate that our JoReS-Diff outperforms state-of-the-art methods on representative benchmarks. We also hope to encourage more works to develop DMs incorporating intrinsic characteristics of low-level vision.
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China under grant U23B2011, 62102069, U20B2063 and 62220106008, the Key R&D Program of Zhejiang under grant 2024SSYS0091, the Sichuan Science and Technology Program under Grant 2024NSFTD0034